Verify: (ASAP) cot(x)=cot(x)-cos(x)/1-sin(x)
Answer by Alan3354(69443)

cot(x)=cot(x)-cos(x)/1-sin(x)
------
cot = cos/sin
cot = (cos/sin - cos)/(1 - sin)
cot = cos*(1/sin - 1)/(1 - sin)
cot = cos*((1 - sin)/sin)/(1 - sin)
Cancel (1 - sin)
cot = cos/sin
cot = cot